 Company just reported Earning Per Share of $50.00 (EPS0=$50)  Management plans to continue to payout 20% of net income in the form of dividends  Before implementing any cost cutting initiatives, management expects earnings will continue to grow at a long term sustainable growth rate of10.0% every year in perpetuity  The company’s cost of equity is 25%

Through cost savings, Company’s management has a plan to increase ROE by 2.5%. Note this will also cause the earnings growth to increase. Recalculate the sustainable long-term growth rate. Using the Dividend Discount Model for constant growth, the value of Company per share is closest to:

a) $86

b) $90

c) $115

d) $215

e) $240
Answer #1

RoE=growth rate/(1-payout rate)=10%/(1-20%)=12.50%
Long term growth rate=(Old RoE+2.5%)*(1-payout rate)=(12.5%+2.5%)*(1-20%)=12%
Value of one share=EPS0*payout rate*(1+g)/(cost of equity-g)=50*20%*(1+12%)/(25%-12%)
=86.1538